# Wake-Informed 3D Path Planning for Autonomous Underwater Vehicles Using A* and Neural Network Approximations

### v1

@@ 8 Conclusions
This study investigated the performance of wake-informed and current-informed path planning approaches for AUVs in complex three-dimensional environments. Both traditional A* algorithms and neural network-based planners were evaluated, with consideration of important metrics such as energy expenditure, path length, encounters with high-velocity cells, turbulent cell count, and computational time. The results demonstrate that the wake-informed A* planner ($W.I._{A*}$) achieves superior energy efficiency and safety by minimizing energy expenditure and avoiding high-velocity regions compared to planners without knowledge of the wake structure ($C.I._{A*}$). $W.I._{A*}$ reduces energy consumption by up to 11.3% compared to the current-informed A* planner, $C.I._{A*}$, at lower speed ranges.
The neural network models ($C.I._{NN}$ and $W.I._{NN}$) offer significant computational speed-ups which makes them highly suitable for real-time applications where computational resources are limited. They are, however, limited in their ability to match the path optimality of the A* planners, exhibiting higher energy expenditures, longer path lengths, and increased encounters with hazardous regions in some cases. While these neural networks provide rapid computations, the approximations may not be sufficient in environments where energy efficiency and safety are paramount. Conversely, the A* planners deliver optimal paths but may be impractical for real-time use due to their computational demands. The wake-informed A* planner’s ability to minimize energy expenditure and avoid high-velocity regions makes it a strong candidate for missions where AUV energy resources are limited and environmental hazards are significant, but where compute remains possible onboard. The slightly increased path lengths are offset by the benefits in energy savings and safety.
The neural network models, while promising for reducing computational overhead, require further refinement to improve their approximation of optimal paths. Enhancements such as incorporating advanced architectures like recurrent neural networks or transformers may help capture the complex spatial-temporal dependencies in three-dimensional environments, not captured by the linear layer structure used here. Expanding the training dataset to include a wider range of scenarios could also improve generalization.
Future research should focus on enhancing the neural network models to better approximate the optimal paths provided by the A* algorithms. Potential avenues include NN architectural improvements, such as exploring more advanced neural network architectures. This could include sequence-to-sequence models or attention mechanisms, to capture complex environmental dynamics. Hybrid Approaches such as the combination of neural networks with traditional algorithms, where the network provides a coarse path that is refined using an optimization algorithm could be investigated to balance speed and optimality. Additionally uncertainty quantification could be investigated to implementing a means to estimate the confidence of neural network predictions, aiding in decision-making processes. By addressing these areas, a better bridge between the performance of neural network-based planners and traditional algorithms could be ensured, enabling efficient and safe real-time path planning for AUVs in complex underwater environments.

### v2

@@ 8 Conclusions
This study demonstrated the substantial benefits of incorporating detailed 3D wake structures into AUV path planning. The wake-informed A* algorithm ($W.I._{A*}$) consistently yielded the most energy-efficient paths, reducing energy consumption by up to 11.3% compared to its current-informed counterpart ($C.I._{A*}$) while also minimizing encounters with high-velocity wake regions. Neural network approximations ($C.I._{NN}$ and $W.I._{NN}$) achieved dramatic computational speed-ups of approximately six orders of magnitude over A*, offering a viable path towards real-time implementation. However, this speed came at the cost of reduced path optimality, with NNs exhibiting 4.5–19.8% higher energy expenditures and generating 9.8–24.4% longer paths. These core findings highlight the critical trade-off between path optimality achievable with detailed environmental information and the computational feasibility required for onboard, real-time AUV operations.
Future research should prioritize bridging the performance gap between the rapid NN approximations and the optimal A* solutions, while also addressing real-world applicability. Key directions include: exploring advanced neural network architectures (e.g., RNNs, Transformers, physics-informed models) better suited for sequential and physics-based tasks; developing hybrid planning systems that leverage NN speed for initial path generation followed by algorithmic refinement; explicitly incorporating AUV dynamic constraints and control limitations into the planning process; and establishing robust methods for uncertainty quantification and safety validation of learning-based planners in complex, dynamic underwater environments. Progress in these areas is essential for enabling safe, efficient, and autonomous AUV operations in close-proximity scenarios characterized by significant hydrodynamic interactions.
Nomenclature
$G$ 3D search grid
$n$ Given node in trajectory
$n^{\prime}$ Neighboring node in trajectory
$N$ Number of nodes
$d$ Euclidean distance
$F_{D}$ Force of drag
$\rho$ Density
$v$ Velocity
$C_{D}$ Coefficient of drag
$A$ Surface area
$E_{D}$ Energy cost over distance
$\omega$ Weighting factor
$c(n)$ Cost
$g(n)$ g-score
$P$ Path of $n$ nodes
$\mathcal{P}$ Set of possible paths
$h(n)$ Search heuristic
$P^{*}_{C}$ Current-informed planner
$P^{*}_{W}$ Wake-informed planner
$E$ Energy (joules)
$L$ Length (meters)
$N_{\text{high-velocity}}(P)$ High velocity cells
$N_{\text{turbulent}}(P)$ Turbulent cells
$\sigma_{v}$ Velocity standard deviation in $G$
$\overline{v}$ Median velocity in $G$
$\delta_{v}$ Indicator function (velocity)
$\delta_{T}$ Indicator function (turbulence)
$\mathcal{L}$ Training loss
$y_{i}$, $\hat{y}_{i}$ True, predicted path coordinates
$m_{i}$ Coordinate path mask
$C.I._{A*}$ Current-informed A* planner
$C.I._{NN}$ Current-informed neural network
$W.I._{A*}$ Wake-informed A* planner
$W.I._{NN}$ Wake-informed neural network
Appendix A Dataset details
The wake flow field data, as discussed in 4.3, was generated from computational fluid dynamics (CFD) calculations undertaken using the commercial software package ANSYS Fluent 2021R1. The flow fields were generated using Reynolds Averaged Navier-Stokes (RANS) simulations with the $k-\epsilon$ turbulence model. An XLUUV geometry based on (Cooper-Baldock et al., 2024) was located in the center of a domain, sized to be 5L (with a characteristic length $L$ of 22 meters) in each axial direction, to ensure boundary effects were mitigated. The $y^{+}$ of the simulation was ensured to be in the range $15030<y^{+}<$ 150, to ensure applicability of the $k-\epsilon$ turbulence model. An INSEAN E1619 propeller model was used, based on existing literature (Cooper-Baldock et al., 2023, 2024). The domain was discretized into 15 million cells, with the majority of these cells clustered in the wake region of the XLUUV using the local refinement method of bodies of influence from (Cooper-Baldock et al., 2024). The freestream velocity was varied from 0.10 m/s to 5.00 m/s in increments of 0.10 m/s and angles of incidence from 0° to 60° degrees of separation in increments of 5°. For each speed and angle pair, a RANS simulation was computed until residual convergence of 1E-5 was achieved. The data from these CFD meshes was then interpolated onto the grid ($G$), using nearest neighbor interpolation from the mesh cells to the cells of the $128^{3}$ grid. The final dataset of distinct grids comprised 500 speed and angle CFD fields. This dataset can be made available upon request.
Appendix B Vehicle details
Table 6 details the AUV model used for the development of the energy cost (Eq. 3). Provided are the physical design it was modeled on, the length, diameter, cross sectional area, drag coefficient and the water density.
Parameter Value Source
AUV Model GRAALTech X300 (Cooper-Baldock et al., 2023; Joung et al., 2012)
AUV Length 3.060 $m$ (Joung et al., 2012)
AUV Diameter 0.254 $m$ (Joung et al., 2012)
AUV Cross sectional area (A) 0.051 $m^{2}$ (Joung et al., 2012)
AUV Drag Coefficient 0.15 (Joung et al., 2012)
Water density 1025.1627 $kg/m^{3}$ (Cooper-Baldock et al., 2024)
Table 6: Physical particulars of the AUV platform used to guide the path planner A* methods and inform the energy compute as listed in Section 4.1
Appendix C NN Implementation
Table 7 provides the general structure, input, layer, output, training design and build of the two NN variants used to replicate the trajectories of the current-informed A* and wake-informed A* planners. These specifications are provided to ensure the work can be replicated. Access to our code repository can be made available upon request.
Table 7: Summary of Neural Network Architecture and Training Configuration.
Category Item Specification
General Network Type Fully Connected Feedforward Neural Network
Purpose Regress sequence of 3D waypoints from input parameters
Variants Two identical networks trained independently ($C.I._{NN}$, $W.I._{NN}$)
Input Dimension 8
Features Start Coords ($x_{s},y_{s},z_{s}$), Goal Coords ($x_{g},y_{g},z_{g}$), Flow Speed ($v$), Flow Angle ($\theta$)
Preprocessing Normalization (Z-score based on training set mean/std, Eq. 16)
Layers Layer 1 (Input) Linear (Dense), Input: 8, Output: 128
Activation 1 ReLU ($\max(0,x)$)
Layer 2 (Hidden) Linear (Dense), Input: 128, Output: 256
Activation 2 ReLU ($\max(0,x)$)
Layer 3 (Output) Linear (Dense), Input: 256, Output: 390
Output Dimension 390
Structure Flattened sequence of 130 waypoints: $(x_{1},y_{1},z_{1},\dots,x_{130},y_{130},z_{130})$
Postprocessing Denormalization (using training set mean/std)
Length Output padded to 130 waypoints; Mask vector used in loss
Training Training Data 14,400 trajectories (80% of A* generated paths)
Validation Data 3,600 trajectories (20% of A* generated paths)
Loss Function Masked Mean Squared Error (MSE, Eq. 17)
Optimizer Adam (Kingma and Ba, 2015)
Learning Rate $1\times 10^{-4}$
Batch Size 64
Max Epochs 50
Initialization Xavier / Glorot (Glorot and Bengio, 2010)
Regularization Early Stopping (Patience: 10 epochs on validation loss)
Build Framework PyTorch (Paszke et al., 2019)